OpenStudy (anonymous):

Some muscle cells have the ability to produce more mitochondria if they are involved in regular exercise. Which answer correctly describes a benefit of this ability?

OpenStudy (anonymous):

Cellular respiration rates would increase in the individual’s cells. The individual would not need as much food as before. More fat molecules would be burned by the individual. The individual’s mass would increase due to the excess mitochondria.

OpenStudy (anonymous):

@Abhisar @aryandecoolest

OpenStudy (abhisar):

What do u think it should be ?

OpenStudy (anonymous):

i think it would be A

OpenStudy (abhisar):

That's \(\huge\checkmark\)

OpenStudy (anonymous):

yeii

OpenStudy (anonymous):

can you help me with one more

OpenStudy (anonymous):

During photosynthesis, unused oxygen atoms form oxygen gas (O2) which is released as a waste product. How did these oxygen atoms originally enter the plant?

OpenStudy (anonymous):

They were part of the ATP molecules gathered by the chloroplasts. They were part of the sugar molecules (C6H12O6) the plant brought in for food. They were part of the carbon dioxide (CO2) molecules that entered through the leaves. They were part of the water molecules (H2O) pulled in through the roots.

OpenStudy (anonymous):

so it would be B?

OpenStudy (abhisar):

This is the reaction of photosynthesis \(\sf 6CO_2+6H_2O--->C_6H_{12}O_6+6O_2\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

si yea, its b lol

OpenStudy (abhisar):

are u sure ?

OpenStudy (anonymous):

or d im no sure

OpenStudy (abhisar):

It's D ;P

OpenStudy (anonymous):

yeiii

OpenStudy (abhisar):

\(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)